what method can i use to solve 2x+3y=9. -3x+3y=-6
Found 2 solutions by Simnepi, lynnlo:
Answer by Simnepi(216)

since both equations have the same number of y's then the best method is elimination.
here's how
2x+3y=9 let's call this equation (1)
-3x+3y=-6 let's call this equation (2)
because they are the same sign, i.e. both positive, then we eliminate them by subtraction. if the signs were different we would eliminate them by adding.
(signs Same, Subtract : signs Different, aDd is how i tell my learners to remember this, emphasising the capital letters!)
now back to the problem
subtracting the equations left hand sides and right hand sides gives
2x+3y-(-3x+3y)=9-(-6)
removing the brackets gives us
2x+3y+3x-3y=9+6
(notice how the negative sign outside the bracket changes all the signs of the terms inside the bracket when it is removed)
simplifying gives
5x=15 so
x=3
now we use this value in one of the original equations i.e.
2x+3y=9
giving
2(3)+3y=9
i.e.
6+3y=9
therefore
3y=3 so
y=1

Answer by lynnlo(4176)

x=3/5
y=13/5
you will use the substition method